# NeRF-To-Real Tester: Neural Radiance Fields as Test Image Generators for Vision of Autonomous Systems

### v1

@@ VI-B RQ2: Diversity
Figure 10: Number of inconsistencies per metric per SUT under different pose transformations and mutations (EF=EfficientNet, MN=MobileNet, VGG=VGG16, XC=Xception, SP=SuperPoint, USP=UnSuperPoint).
(a) Cos. Similarity
(b) L2 Norm
(c) Repeatability
(d) IP Spread
Figure 11: Number of inconsistencies per metric per scene under different pose transformations and mutations.
dory1 dory2 dory3 dory4 dory5
#Test Images 1806 1980 2148 1818 2136
plane dozer car truck bike
#Test Images 1902 2154 2286 2508 1818
TABLE IV: Number of NeRF-generated test images per scene.
@@ VI-C RQ3: Efficiency
All NeRF trainings and image renderings are performed on two NVIDIA GPUs of type RTX 3090 Ti. Since the training and rendering times depend on the type of hardware available, we report carbon emissions to enable performance comparison. Once we have the fully trained models and rendered images of the scenes, the efficiency of executing N2R-Tester is predominantly determined by the efficiency of the SUTs, i.e. the time taken to process a test image, thus we consider it most appropriate to report the time taken for N2R-Tester to generate such test images.
NeRF Training Times $\downarrow$
See Tbl. I for a summary of the training times of 30k steps for each of the NeRF models.
NeRF Rendering Frames per Second (FPS) $\uparrow$
We report the average FPS for a trained NeRF model to render images at various image resolutions in Tbl. V. DeepXplore mutations can be artificially engineered with OpenCV at around 300 FPS. Compared to engineering a 3D environment in simulation, which requires hiring an experienced software engineer and many working hours, NeRF models still offer an effective shortcut to realistic 3D scene reconstruction.
@@
TABLE V: Rendering frame rates for a trained NeRF model at different image resolutions in frames per second (FPS) $\uparrow$.
Carbon Emissions $\downarrow$
An estimate of 300 hours of GPU time for NeRF training and 200 hours for image rendering were required on private infrastructure, having a carbon efficiency of 0.432 kgCO2eq/kWh [38]. For our type of GPUs the total emissions are estimated to be 75.6 kgCO2eq by the Machine Learning Impact calculator, an equivalent of 305km driven by an average combustion-engine passenger car. Lower carbon emissions mean better performance. Notably, the vast majority of parameter fine-tuning was required for the underwater models. We suspect the challenge was aggravated by colmap’s difficulty to find coherent camera poses when faced with underwater images containing moving elements and backscattering light effects in the scenes. We consider it worth investigating if other NeRF architectures could mitigate these effects.
@@ VI-D Threats to Validity and Limitations
The test results with NeRF-generated data for IPDs and classifiers do not necessarily transfer to other scenes, other IPDs, other classification domains, and other classification models. This is an inherent problem with machine language models. So far the only way to increase confidence of transferability is to accumulate experience with more domains, more models, and more vision components. In our experiments, we have used two vastly different visual tasks (classification and interest-point detection) and used data sets from different domains and different locations. We have used both DNN-based and non-DNN-based IPDs, and large and small classifier models, to increase generalisability.
@@
Similarly to any DNN, the output quality during inference relies heavily on the relevant samples received during training. For NeRFs, the visual fidelity in the generated images depends on the number of ray samples in proximity. Consequently, the images of the original camera poses $(\mathbf{x},\mathbf{d})$ of the real images in the scene $i_{\textrm{real}}\!\in\!S$ are most likely to be most accurately represented in the NeRF model. Conversely, the further we deviate from these poses, the likelihood of rendering under-defined pixels increases.
Figure 12: Examples of NeRF artifacts (enlarged, best seen on screen). Top left: a real image of a reflective surface. Top right: an imprecise rendering in car of the same view as top left. Bottom left: An example of an under-defined pixel patch in bike. Bottom right: An example of a ghost wall in dory1.
@@ VII Related Work
Since DNNs are efficient in processing image data, there is a trend to replace traditional vision-based navigational algorithms by end-to-end trained neural networks. It is thus crucial to consider the challenges associated with testing neural networks in general. Methods such as DeepXplore [6] and DeepTest [7] attempt to uncover erroneous behaviour of neural networks by synthesising inputs based on maximising neuron coverage. However, while doing this, they do not respect the geometric properties of the scene, so the generated sequences of images do not guarantee spatio-temporal consistency. DeepRoad is a MT framework producing test images with different weather effects via domain adaptation with GANs [12]. The images are transformed on a per-frame basis, so DeepRoad is limited to test of aspects such as steering angle accuracy. It does not offer geometric consistency for a stream of frames, hence it is unsuitable for testing vSLAM components such as IPDs processing consecutive frames. MT has been applied to evaluate aerial drone control policies in simulation [39]. Here, the SUT is the drone controller and test inputs are various states of the simulated environment in combination with a mission, e.g. ‘take-off’ or ‘fly-to-location’ This line of work was extended to develop MRs for testing obstacle-avoidance algorithms for single and multiple agents in simulation [40]. Our work addresses testing image processing algorithms with images such as test inputs, hence posing as a distinct approach to applying metamorphic testing in the robotics domain. To the best of our knowledge, we are the first to investigate how NeRFs can be applied in test image generation.
NeRFs for localisation, navigation and scene representation
The potential applications of NeRFs in the robotics domain are still largely unexplored [41]. NeRFs can improve vision-based trajectory planners and collision avoidance for aerial drones [42]. NeRF2Real fuses a NeRF with a physics simulator to achieve fast renderings of a scene as well as modeling of dynamic objects, the robot body, interactions, and collisions [43]. SPARTN is a data augmentation technique for an eye-in-hand robotic arm that feeds NeRF-generated visual samples of grasping behaviours into a robotic imitation learning pipeline [44]. Notably, SPARTN and NeRF2Real conduct evaluations on physical robots and hence directly tackle the “simulation-to-reality” gap. While these works enhance robotics functionality, we focus on automating meaningful test data generation for systematic vision task testing in robotics. Several works address the degradation of NeRFs visual clarity due to inaccurate or inaccessible camera poses during training. Bundle-Adjusting Neural Radiance Fields [45] optimize jointly for camera poses and visual input, therefore alleviating the dependence on the accuracy of Structure-from-Motion tools, and enabling reconstruction from imperfect or even unknown camera poses [17]. Renderable Neural Radiance Maps enhance NeRFs’ ability to map visual input into latent spaces, providing additional 3D scene information for improved localization [46]. Loc-NeRF combines Monte-Carlo localisation with a pre-trained NeRF for fast localisation without depending on a perfect initial pose estimate [47]. These methods improve visual pose estimation and navigation through scenes, distinct from independently building realistic test mechanisms to uncover faulty behaviours of navigation components, as presented in this paper. NeRF-SLAM uses the output of dense monocular SLAM, i.e. depth maps and camera poses with the associated uncertainties for NeRF training [48]. NeRF-SLAM addresses neither testing nor improve SLAM with NeRFs, as the name may suggest. It aims to produce renderings with high photometric and geometric accuracy. SeaThru-NeRFs improve generated images by differentiating back-scatter effects from large unbounded underwater scenes [49]. In this work, we used small and bounded underwater scenes where these advantages did not offer significant visual improvements. Beyond-NeRF Underwater targets colour correction of images captured by a bottom-facing AUV camera [50]. These typically feature an artificial light source causing a high concentration of unexposed pixels, which makes them unsuitable in our context. Finally, we were unable to find and access source code for WaterNeRF [51] and WaterHE-NeRF [52].
@@ VIII Conclusion
N2R-Tester is a metamorphic testing and test data synthesis method for perception components of autonomous vehicles. Like other metamorphic methods, it does not use domain-specific or project-specific rules to specify tests. N2R-Tester offers a plethora of advantages over other image generation techniques for navigation applications. First, it grants precise control over image transformations, as opposed to models such as GANs. Self-contained image rotations are fundamentally not possible using real images alone, without loss of information; as soon as we rotate an image for example in OpenCV, we loose pixels in the image corners and need to zoom in, extrapolate, or inpaint to get a coherent image. Second, unlike other computer graphics methods, such as textured meshes combined with sophisticated rendering engines, it does not require explicit 3D models of the scene. A NeRF model is a continuous function approximation capable of implicitly representing complex scenes with arbitrary shapes, appearances and textures, which offers great flexibility over traditional techniques. Third, it does not require human annotation. A NeRF model learns to represent a 3D scene solely based on real 2D image inputs. Fourth, N2R-Tester does not directly rely on image fidelity metrics such as PSNR, SSIM, and LPIPS for assessing image realness on a per-sample basis, rather they act as more global indicators of satisfactory NeRF in-training-evaluation. N2R-Tester can be potentially extended to test stereo-vision, as it can render two offset camera views. NeRFs inherently model depth in the $\alpha$-channel of the output layer and could therefore also be used to test depth perception. Finally, NeRFs could be replaced by newer 3D reconstruction models such as Gaussian Splatting [53].

### v2

@@ VI-B RQ2: Diversity
Figure 9: Number of inconsistencies per metric under N2R transformations (red) and DeepXplore mutations (blue).
In Fig. 9 we show the number of inconsistencies across all SUTs according to our MT framework with three different threshold values for $\epsilon=[0.1,0.2,0.5]$. This means that we consider a deviation of a SUT metric by more than 10%, 20%, 50% an inconsistency in the behavior of a SUT. The strictest test is $\epsilon=0.1$. N2R manages to trigger a higher median of inconsistencies for L2 Norm and IP Spread than DeepXplore mutations, as well as better performance on the strictest tests with $\epsilon=0.1$ according to the repeatability metric. See Fig. 10 for two examples of inconsistent SUT behavior.
Figure 10: Examples of inconsistent SUT behaviors: Image (a) transformed by $\tau_{4}$ is classified as ImageNet1k class label “trash can” by MobileNet with 0.53 confidence while having average image metric values (PSNR 15.77, SSIM 0.45, LPIPS 0.72). Image (b) transformed by $\tau_{4}$ has a low repeatability score of 0.37 on SuperPoint. The low score is caused by a low agreement of points from the associated real image (green) compared to points found in this image (red). Image (b) has good values for PSNR 35.12, SSIM 0.70, LPIPS 0.60.
@@ VI-C RQ3: Efficiency
All NeRF trainings and image renderings are performed on two NVIDIA GPUs of type RTX 3090 Ti, averaging $\sim 1.5$ hours of training per model (see Tbl. I). Once we have all rendered images of the scenes, the efficiency of executing N2R-Tester is determined by the efficiency of the SUTs processing times. Thus we consider it most appropriate to report the time taken for N2R-Tester to generate test images.
We report the average FPS for a trained NeRF model to render images at various image resolutions in Tbl. II. DeepXplore mutations can be artificially engineered with OpenCV at around 300 FPS. NeRF models still offer an effective shortcut to realistic 3D scene reconstruction compared to simulating a 3D environment, which requires hiring an experienced software engineer and many working hours.
@@
TABLE II: Rendering frame rates for a trained NeRF model at different image resolutions in frames per second (FPS) $\uparrow$.
@@ VI-D Threats to Validity and Limitations
Figure 11: Left: An example of an under-defined pixel patch in bike. Right: An example of a ghost wall in dory1.
@@
The test results with NeRF-generated data for IPDs and classifiers do not necessarily transfer to other scenes, other IPDs, other classification domains, and other classification models. So far the only way to increase confidence of transferability is to accumulate experience with more domains, more models, and more vision components. In our experiments, we have used two vastly different visual tasks (classification and interest-point detection) and used data sets from different domains and different locations. We have used both DNN-based and non-DNN-based IPDs, and large and small classifier models, to increase generalization.
The NeRF reconstruction process is initialized with random pixels, which gradually converge to a geometrically consistent scene during training. During inference this presents an advantage; NeRFs fail gracefully with ambiguity in the reconstruction, when rendering an image at a view with insufficient radiance samples. This can result in “artifacts” or “ghost walls”, i.e. image patches appearing as random noise or blurry (Fig. 11). Researchers in the NeRF community are in a race trying to outperform each other in mitigating these pixel effects. We are curious as to how exactly these phenomena potentially affect the performance of the SUTs, however this exploration marks a whole line of research in its own right.
@@ VII Related Work
DeepXplore [5], DeepTest [6] and DeepRoad [9] aim to uncover erroneous behavior of neural networks by synthesizing inputs on a per-frame basis. However they cannot keep geometric properties of a 3D scene or guarantee spatio-temporal consistency. Hence they are unsuitable for testing vSLAM components such as IPDs processing consecutive frames. MT has been applied to evaluate aerial drone control policies in simulation [22], [23]. Here, the SUT is the drone controller and test inputs are various states of the simulated environment in combination with a mission. Our work addresses testing image processing algorithms with images as test inputs, hence posing as a distinct approach to applying metamorphic testing in the robotics domain.
NeRFs in Robotics
NeRFs can improve vision-based trajectory planners and collision avoidance for aerial drones [24]. NeRF2Real fuses a NeRF with a physics simulator to achieve fast renderings of a scene and modeling of dynamic objects, the robot body, interactions, and collisions [25]. SPARTN is a data augmentation technique for an eye-in-hand robotic arm that feeds NeRF-generated visual samples of grasping behaviors into an imitation learning pipeline [26]. These works improve robotics functionality, however we automate test data generation for vision tasks in robotics.
NeRFs for high quality rendering
Several works aim to improve of NeRFs visual clarity caused by inaccurate or inaccessible camera poses during training, e.g. Bundle-Adjusting Neural Radiance Fields [27] or Loc-NeRF [28]. These methods improve visual pose estimation and navigation through scenes. We use NeRFs as realistic test mechanisms to uncover faulty behaviors of navigation components. Other types of NeRF architectures reconstruct 3D underwater scenes by reducing back-scatter effects [29] or color distortions [30].
@@ VIII Conclusion
N2R-Tester is a metamorphic testing and test data synthesis method for perception components of autonomous vehicles. Like other metamorphic methods, it does not use domain-specific or project-specific rules to specify tests. N2R-Tester offers a plethora of advantages over other image generation techniques for navigation applications. First, it grants precise control over image transformations, as opposed to models such as GANs. Second, unlike other computer graphics methods, such as textured meshes combined with sophisticated rendering engines, it does not require explicit 3D models of the scene. A NeRF model is a continuous function approximation capable of implicitly representing complex scenes with arbitrary shapes, appearances and textures, which offers great flexibility over traditional techniques. Third, it does not require human annotation. Finally, N2R-Tester can be potentially extended to test stereo-vision, as it can render two offset camera views. NeRFs inherently model depth in the $\alpha$-channel of the output layer and could therefore also be used to test depth perception.